What is the value of x to the 2nd power equals 36/81 explain?
Answer by Alan3354(69443)

What is the value of x to the 2nd power equals 36/81 explain?
---------
x^2 = 36/81 = 4/9
----
(2/3)^2 = 4/9
(-2/3)^2 = 4/9